OpenStudy (anonymous):

Locate the absolute extrema of the function on the closed interval: y = 3x^(2/3) - 2x, [-1, 1]

OpenStudy (anonymous):

\[y' = \frac{2}{x^{1/3}} - 2\] \[y' = 2(\frac{1-x^{1/3}}{x^{1/3}})\] \[2(\frac{1-x^{1/3}}{x^{1/3}})=0\] \[\frac{1-x^{1/3}}{x^{1/3}} = 0\]

OpenStudy (anonymous):

^-- finding critical points, btw

OpenStudy (anonymous):

\[1-x^{1/3} = 0\] \[x^{1/3} = 1\] \[x = 1\]?? my book reads "0" as the critical point.

OpenStudy (anonymous):

0 is a critical point because that is where the derivative is undefined. the definition of critical point is where f' = 0 OR f' does not exist.

OpenStudy (anonymous):

i understand, but by setting the derivative = to 0 and solving for x, i've defined the critical point for which the derivative of the function is equal to zero, no? as shown in my work above

OpenStudy (anonymous):

yes... so in essence, in the interval [-1, 1], you have two critical numbers.

OpenStudy (anonymous):

ummm wait... are you looking at [-1, 1] or (-1, 1) ???

OpenStudy (anonymous):

in the closed interval [-1,1]